# Dyn-HTE: High-temperature expansion of the dynamic Matsubara spin correlator

### v1

@@ Abstract
The high-temperature series expansion for quantum spin models is a well-established tool to compute thermodynamic quantities and equal-time spin correlations, in particular for frustrated interactions. We extend the scope of this expansion to the dynamic Matsubara spin-spin correlator and develop a fully analytic algorithm to compute its expansion coefficients. We focus on Heisenberg models with a single coupling constant $J$ and spin lengths $S\in\{1/2,1\}$. The expansion coefficients up to 12th order in $J/T$ are precomputed on all possible $\sim 10^{6}$ graphs embeddable in arbitrary lattices and are provided under https://github.com/bsbierski/Dyn-HTE. This enables calculation of static momentum-resolved susceptibilities for arbitrary site-pairs or wavevectors. We test our results for the $S=1/2$ Heisenberg chain and on the triangular lattice model. Moreover, the analytic frequency dependence in the expansion allows for stable analytic continuation to the real-frequency dynamic structure factor. This important application is discussed in a companion letter.
@@ I Introduction
The high-temperature series expansion (HTE) is an invaluable tool for the theoretical analysis of quantum spin systems in thermal equilibrium, for monographs see, e.g., Refs. 1, 2. The HTE is oblivious to frustration and entanglement, is formulated directly in the thermodynamic limit and remains applicable in high dimensions. Up to now, the HTE can target thermodynamic quantities like entropy, heat capacity or uniform susceptibility and was also applied to obtain equal-time spin correlation functions $G_{ii^{\prime}}^{zz}=\left\langle S_{i}^{z}S_{i^{\prime}}^{z}\right\rangle$ for arbitrary site-pairs $ii^{\prime}$. Key technical advancements of the HTE method included the extension to arbitrary spin-length $S$ [3, 4], open-source software packages [5] or the inclusion of magnetic fields [6]. As the name suggests, HTE is mainly challenged at temperatures much lower than the typical spin interaction $T\ll J$ when resummation schemes of the bare series give ambiguous and thus non-reliable results. To some extent, these issues can be bypassed if qualitative information on the very low-$T$ behavior is available [7].
@@
However, in order to achieve a better understanding of collective phenomena arising in equilibrium quantum spin systems it is mandatory to also computationally target spin dynamics. Here, the most elementary observable is the dynamical spin-spin correlator $G^{zz}_{ii^{\prime}}(t)=\left\langle S_{i}^{z}(t)S_{i^{\prime}}^{z}\right\rangle$ with $S_{i}^{z}(t)=e^{iHt}S_{i}^{z}e^{-iHt}$ representing a local spin operator in the Heisenberg picture [12]. The dynamical structure factor which is defined as a spatial and temporal Fourier transform of $G^{zz}_{ii^{\prime}}(t)$ contains rich information on the dipolar excitation spectrum, the presence and stability of quasiparticles or the fractionalization of spin-flip excitations [13]. It is routinely measured using inelastic neutron scattering [14] in the solid-state context and recently also for cold-atom quantum simulators [15] via Raman spectroscopy. Although the dynamical structure factor can be calculated in various ways for different settings [16] (e.g. via exact diagonalization and linked-cluster methods [17], spin-wave theory [18] and tensor-networks [19]), it is still desirable to approach spin dynamics via the HTE and benefit from its above mentioned strengths. This is particular important for (strongly frustrated) spin liquid candidates [20] where the observables available from conventional HTE are often quite featureless as functions of temperature and momentum. A natural intermediate goal to this endeavor is the extension of the HTE for the dynamical spin correlation function in imaginary time $t\rightarrow-i\tau\in i\mathbb{R}$ which is a somewhat simpler task given the benefits of the Matsubara formalism for perturbation theory in equilibrium [12].
@@
In this work we develop this extension of the HTE to the dynamic imaginary frequency (Matsubara) spin-spin correlator. We term this method dynamic HTE (Dyn-HTE). We lay out the general formalism and apply it to Heisenberg models with a single coupling constant $J$ and spin lengths $S\in\{1/2,1\}$. The coefficients for the expansion in powers of $x\equiv J/T$ (and in inverse frequency) up to order $n_{max}=12$ are precomputed exactly on all possible $\sim 10^{6}$ graphs and are offered for download along with powerful tools for the creation of arbitrary lattices and efficient graph embeddings [21]. On the technical side Dyn-HTE hinges on the exploitation of the recently developed Kernel trick [22] which solves the $(n+2)$-fold imaginary-time integrals (for expansion order $n$) analytically. Interestingly, the numerical cost for the expansion in order $n$ is only modestly increased compared to the conventional HTE for the equal-time correlator, depending on graph topology. In the worst case, the extra effort is by a factor of $n$. The high expansion orders set Dyn-HTE apart from recently developed alternative spin-diagrammatic expansions of the 1-J-irreducible part of the Matsubara correlator [23, 24] which are built on different diagrammatic schemes [25, 26, 27, 28, 29].
One experimental observable directly available from Dyn-HTE is the static susceptibility for arbitrary site-pairs. We consider the AFM $S=1/2$ Heisenberg chain and compare the static susceptibility from Dyn-HTE against error controlled quantum Monte Carlo (QMC). We also report the static susceptibility for the frustrated triangular lattice model. Finally, the Matsubara correlator is related to the real-frequency dynamical spin structure factor by analytic continuation [12]. Despite being an ill-conditioned and error-prone procedure for numerical data generated with most other imaginary-time approaches [30, 31, 32, 33], the analytic frequency dependence maintained in Dyn-HTE is instrumental to bypass these difficulties and obtain stable results. This is described in our companion work [34] along with benchmark examples and numerous applications.
@@ II Heisenberg spin model
We consider a system of length-$S$ quantum spins with operators $S^{\alpha}_{i}$ where $\alpha=x,y,z$. The subscript $i=1,2,...,N$ refers to the site at position $\mathbf{r}_{i}$ of an arbitrary lattice $\mathcal{L}$. The spins interact via Heisenberg exchange characterized by a single coupling constant $J$ along an arbitrary subset of all $N(N-1)/2$ site pairs which we call bonds $(ii^{\prime})$. This includes the important case of symmetry-related nearest-neighbor interactions, but also all-to-all or spatially disordered (but equal) interactions. The Hamiltonian reads
@@ III Matsubara spin correlator and Dyn-HTE
where $\nu_{m}=2\pi Tm$ with $m\in\mathbb{Z}$ is a (bosonic) Matsubara frequency and $S_{i}^{z}(\tau)=e^{H\tau}S_{i}^{z}e^{-H\tau}$ denotes imaginary time evolution. The operator $\mathcal{T}$ enforces imaginary time ordering, the operator with larger imaginary time argument is moved to the left. Note that the double integral is often replaced by a single integral over the difference $\tau-\tau^{\prime}$ but the form in (4) will prove useful later. Due to spin rotation symmetry in the Heisenberg Hamiltonian, Eq. (4) also determines correlations for all other spin-flavor combinations, $G^{\alpha\alpha^{\prime}}=\delta_{\alpha\alpha^{\prime}}G^{zz}$ and the $zz$ superscript of $G^{zz}$ is dropped in the following to ease the notation. The spatial Fourier transform of Eq. (4) reads
@@
As we show in Sec. IV, the Dyn-HTE of the Matsubara correlator (4) is in fact a double expansion in $x=J/T$ and inverse Matsubara frequency $1/\nu_{m}$. To order $n_{max}$ in $x$, Dyn-HTE reads
@@
$\begin{cases}p_{ii^{\prime}}^{(0)}(x)\!\!&\!:\!m=0\\ \sum_{r=1}^{r_{max}}p_{ii^{\prime}}^{(2r)}(x)\left(\frac{x}{2\pi m}\right)^{2r}\!\!&\!:\!m\neq 0\end{cases}\!+O(x^{n_{max}+1})$ (7)
where $r_{max}=\left\lfloor n_{max}/2\right\rfloor$ is the integer floor of $n_{max}/2$. The dimensionless polynomials $p_{ii^{\prime}}^{(2r)}(x)$ are of degree $n_{max}-2r$ with (real) rational coefficients. Depending on the lattice and site-pair $ii^{\prime}$, some of the polynomial coefficients can be zero. Note that the form (7) is consistent with the symmetries of the Matsubara correlator discussed above.
Our open-source numerical implementation [21] currently provides the polynomial coefficients in Eq. (7) exactly for arbitrary lattices $\mathcal{L}$, all site-pairs $ii^{\prime}$ therein and for $S\in\{1/2,1\}$ and $n_{max}=12$.
@@
At finite frequency $m\!\neq\!0$, the Matsubara correlator does not directly represent an observable physical quantity. It is mainly considered for the simplicity of the diagrammatic field-theoretical framework which can be employed for its computation [12]. The real-time spin-spin correlator $G_{ii^{\prime}}(t)=\left\langle S_{i}^{z}(t)S_{i^{\prime}}^{z}\right\rangle$ or rather its temporal and spatial Fourier transform, the dynamical structure factor $S(\mathbf{k},\omega)$, relates to $G_{\mathbf{k}}(i\nu_{m})$ via analytical continuation,
$G_{\mathbf{k}}(i\nu_{m})=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}\omega\,\frac{A_{\mathbf{k}}(\omega)}{\omega-i\nu_{m}}.$ (9)
and the fluctuation-dissipation relation $A_{\mathbf{k}}(\omega)=2\pi(1-e^{-\omega\beta})S(\mathbf{k},\omega)$ which connects the imaginary part of the dynamical spin susceptibility $A_{\mathbf{k}}(\omega)$ (also known as the spectral density) to the dynamical structure factor [12]. Importantly, Eq. (9) is strictly correct only for $i\nu_{m}\neq 0$ and an additional term might appear for $i\nu_{m}=0$ [35, 36, 37]. This correction quantifies a long-term memory effect for spin correlations which is a measure of non-ergodicity. It was also shown to quantify the difference between the static (isothermal) susceptibility and the zero-frequency limit of the isolated Kubo susceptibility $-G_{\mathbf{k}}^{R}(\omega)$.
There is no shortage in established methods to calculate the numerical value of the Matsubara spin correlator $G_{\mathbf{k}}(i\nu_{m})$ at finite temperature and for some given frequencies $i\nu_{m}$ for relevant models, most notably the error-controlled quantum Monte Carlo [30] (for sign-problem free models) or diagrammatic approaches like pseudo-fermionic functional renormalization [38, 32, 33] or the bold-line diagrammatic Monte Carlo technique [31]. Consequently, a large body of work has been dedicated on the notoriously hard analytical continuation problem [39, 40, 41, 42, 43], which is inverting Eq. (9). What separates our Dyn-HTE (7) from these approaches is that it provides the analytic dependence of $G_{\mathbf{k}}(i\nu_{m})$ in (inverse) frequency. In our companion paper [34], we show that this allows for a different and quite stable analytical continuation procedure based on the concept of frequency moments and continued fraction expansions of the spectral density [44, 45, 46, 47]. There, we also discuss benchmarks and applications all centered around the real-frequency dynamical structure factor. In contrast, the remainder of the current work forms the basis of this application in that it will be concerned with the efficient evaluation of the Dyn-HTE, i.e. $p_{ii^{\prime}}^{(2r)}(x)$ from Eq. (7). We also present, benchmark and discuss results for the Matsubara correlator in imaginary frequency.
@@ IV Derivation of Dyn-HTE
In this central technical section we derive the Dyn-HTE as in Eq. (7) and provide an efficient algorithm its calculation. To start, we split the Hamiltonian (1) into a non-interacting part $H_{0}=0$ (vanishing in the absence of a magnetic field) and an interacting part $V$ (here the full $H$) and define the dimensionless expansion coefficients $c_{ii^{\prime}}^{(n)}(i\nu_{m})$ via
@@
The lowest-order $n=0$ term represents the Curie susceptibility of a free spin which is local and static (has only a contribution at $m=0$), $c_{ii^{\prime}}^{(0)}(i\nu_{m})=\delta_{ii^{\prime}}\delta_{0,m}S(S+1)/3$.
@@ IV.1 Expansion coefficients
General perturbation theory for the Matsubara correlator [12] provides the remaining expansion coefficients at order $n=1,2,...$ as the expression
$c_{ii^{\prime}}^{(n)}(i\nu_{m})=\frac{T^{n+1}}{n!}\sum_{b_{1},...,b_{n}}\int_{0}^{\beta}e^{i\nu_{m}(\tau-\tau^{\prime})}\mathrm{d}\tau_{1}...\mathrm{d}\tau_{n}\mathrm{d}\tau\mathrm{d}\tau^{\prime}\left\langle\mathcal{T}V_{b_{1}}(\tau_{1})...V_{b_{n}}(\tau_{n})S_{i}^{z}(\tau)S_{i^{\prime}}^{z}(\tau^{\prime})\right\rangle_{0}^{\mathrm{V-con}}.$ (11)
Here the sum is over bonds $b_{l}=(i_{l}i_{l}^{\prime})$. The operator time evolutions $V_{b_{1}}(\tau_{1})...V_{b_{n}}(\tau_{n})S_{i}^{z}(\tau)S_{i^{\prime}}^{z}(\tau^{\prime})$ are now with respect to $H_{0}$ only and thus trivial for our case, $H_{0}=0$, e.g. $V_{b_{1}}(\tau_{1})=V_{b_{1}}$. Nevertheless we must keep the time arguments because by virtue of $\mathcal{T}$ they determine the ordering of the string of generally non-commuting operators. These operators are thus distinguishable regardless of the chosen bonds $b_{l}$. The $\left\langle...\right\rangle_{0}$ denotes a thermal average with respect to Hamiltonian $H_{0}$. Since $H_{0}=0$, the former is simply an (effectively) infinite temperature average or normalized trace over the whole $N$-site spin Hilbert-space, $\left\langle...\right\rangle_{0}=\mathrm{tr}[...]/(2S+1)^{N}$ which factorizes according to site-index, e.g. $\left\langle S_{1}^{z}S_{2}^{+}S_{1}^{z}S_{2}^{-}\right\rangle_{0}=\left\langle S_{1}^{z}S_{1}^{z}\right\rangle_{0}\left\langle S_{2}^{+}S_{2}^{-}\right\rangle_{0}$ with the on-site operator order maintained.
The superscript ($\mathrm{V-con}$) on the right-hand side of Eq. (11) refers to the $V$-connected correlator with $n$ insertions of $V$ which is of order $V^{n}\sim J^{n}$ and contains all contributions from the expansion of the partition function $Z$ in the denominator on the right-hand side of Eq. (2). Following Ref. 48, it can be written via recursive subtractions from the ”full” correlator (without $\mathrm{V-con}$)
@@
-- $\!\!\sum_{S\subsetneq\{1,...,n\}}\left\langle\mathcal{T}\left[\prod_{k\in S}V_{b_{k}}(\tau_{k})\right]S_{i}^{z}(\tau)S_{i^{\prime}}^{z}(\tau^{\prime})\right\rangle_{0}^{\mathrm{V-con}}\left\langle\mathcal{T}\prod_{l\in\{1,..,n\}\backslash S}V_{b_{l}}(\tau_{l})\right\rangle_{0},$
@@
$c_{ii^{\prime}}^{(n)}(i\nu_{m})$ $=$ $\sum_{b_{1},...,b_{n}}\frac{T^{n+1}}{n!}\int_{0}^{\beta}e^{i\nu_{m}(\tau-\tau^{\prime})}\mathrm{d}\tau_{1}...\mathrm{d}\tau_{n}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
$\times$ $\{\left\langle\mathcal{T}V_{b_{1}}(\tau_{1})...V_{b_{n}}(\tau_{n})S_{i}^{z}(\tau)S_{i^{\prime}}^{z}(\tau^{\prime})\right\rangle_{0}-\!\!\!\!\!\sum_{S\subsetneq\{1,...,n\}}\!\!\left\langle\mathcal{T}\!\!\left[\prod_{k\in S}V_{b_{k}}(\tau_{k})\right]\!S_{i}^{z}(\tau)S_{i^{\prime}}^{z}(\tau^{\prime})\!\!\right\rangle_{0}^{\!\!\mathrm{V-con}}\!\!\left\langle\mathcal{T}\!\!\!\!\!\prod_{l\in\{1,..,n\}\backslash S}\!\!\!\!\!V_{b_{l}}(\tau_{l})\right\rangle_{\!\!0}\!\}$
$\equiv$ $\sum_{b_{1},...,b_{n}}F^{(n)}_{ii^{\prime}}\left(b_{1},...,b_{n};i\nu_{m}\right).$ (14)
The last line explicitly exposes the spatial structure of the different terms in the sum using bonds $b_{1},...,b_{n}$ which are not necessarily all distinct.

### v2

# High-temperature series expansion of the dynamic Matsubara spin correlator

@@ Abstract
The high-temperature series expansion for quantum spin models is a well-established tool to compute thermodynamic quantities and equal-time spin correlations, in particular for frustrated interactions. We extend the scope of this expansion to the dynamic Matsubara spin-spin correlator and develop an algorithm that yields exact expansion coefficients in the form of rational numbers. We focus on Heisenberg models with a single coupling constant $J$ and spin lengths $S\in\{1/2,1\}$. The expansion coefficients up to 12th order in $J/T$ are precomputed on all possible $\sim 10^{6}$ graphs embeddable in arbitrary lattices and are provided under https://github.com/bsbierski/Dyn-HTE. This enables calculation of static momentum-resolved susceptibilities for arbitrary site-pairs or wavevectors. We test our results for the $S=1/2$ Heisenberg chain and on the triangular lattice model. An important application that we discuss in a companion letter is the calculation of real-frequency dynamic structure factors. This is achieved by identifying the high-frequency expansion coefficients of the Matsubara correlator with frequency moments of the spectral function.
@@ I Introduction
The high-temperature series expansion (HTE) is an invaluable tool for the theoretical analysis of quantum spin systems in thermal equilibrium, for monographs see, e.g., Refs. 1, 2. The HTE is oblivious to frustration and entanglement, is formulated directly in the thermodynamic limit and remains applicable in high dimensions. Up to now, the HTE can target thermodynamic quantities like entropy, heat capacity or uniform susceptibility and was also applied to obtain equal-time spin correlation functions $G_{ii^{\prime}}^{zz}=\left\langle S_{i}^{z}S_{i^{\prime}}^{z}\right\rangle$ for arbitrary site-pairs $ii^{\prime}$. Key technical advancements of the HTE method included an extension to arbitrary spin-length $S$ [3, 4], flexible open-source software packages [5] or inclusion of magnetic fields [6]. As the name suggests, the main challenge of HTE is to reach temperatures much below the spin interaction $T\ll J$ when resummation schemes of the bare series give ambiguous and thus non-reliable results. To some extent, these issues can be bypassed if qualitative information on the very low-$T$ behavior is available [7].
@@
However, in order to achieve a better understanding of collective phenomena arising in equilibrium quantum spin systems it is mandatory to also computationally target spin dynamics. Here, the most elementary observable is the dynamical spin-spin correlator $G^{zz}_{ii^{\prime}}(t)=\left\langle S_{i}^{z}(t)S_{i^{\prime}}^{z}\right\rangle$ with $S_{i}^{z}(t)=e^{iHt}S_{i}^{z}e^{-iHt}$ representing a local spin operator in the Heisenberg picture [12]. The dynamical structure factor which is defined as a spatial and temporal Fourier transform of $G^{zz}_{ii^{\prime}}(t)$ contains rich information on the dipolar excitation spectrum, the presence and stability of quasiparticles or the fractionalization of spin-flip excitations [13]. It is routinely measured using inelastic neutron scattering [14] in the solid-state context and recently also for cold-atom quantum simulators [15] via Raman spectroscopy. Although the dynamical structure factor can be calculated in various ways for different settings [16] (e.g. via exact diagonalization and linked-cluster methods [17], spin-wave theory [18] and tensor-networks [19]), it is still desirable to approach spin dynamics via the HTE and benefit from its above mentioned strengths. This is particular important for (strongly frustrated) spin liquid candidates [20] where the observables available from conventional HTE are often featureless as functions of temperature and momentum. A natural intermediate goal to this endeavor is the extension of the HTE for the dynamical spin correlation function in imaginary time $t\rightarrow-i\tau\in i\mathbb{R}$ which is a somewhat simpler task given the benefits of the Matsubara formalism for perturbation theory in equilibrium [12].
@@
In this work, we develop this extension of the HTE to the dynamic imaginary frequency (Matsubara) spin-spin correlator. We term this method dynamic HTE (Dyn-HTE). We lay out the general formalism and apply it to Heisenberg models with a single coupling constant $J$ and spin lengths $S\in\{1/2,1\}$. The coefficients for the expansion in powers of $x\equiv J/T$ up to order $n_{max}=12$ (and in inverse frequency to order $\left\lfloor n_{max}/2\right\rfloor$) are precomputed exactly in the form of rational numbers on all possible $\sim 10^{6}$ graphs and are offered for download along with powerful tools for the creation of arbitrary lattices and efficient graph embeddings [21]. On the technical side Dyn-HTE hinges on the exploitation of the recently developed Kernel trick [22] which solves the $(n+2)$-fold imaginary-time integrals (required at expansion order $n$) analytically. Interestingly, the numerical cost for the expansion in order $n$ is only modestly increased compared to the conventional HTE for the equal-time correlator. Depending on graph topology the extra effort is by a factor of $n$ in the worst case.
For historic context, early developments of spin-diagrammatic schemes date back to the late 1960s [23, 24] where the Wick theorem was generalized for spin operators, see also Ref. [25] for a modern recursive formulation. Recently, these ideas have partially been revived in the development of a functional-renormalization group approach for spin systems [26, 27, 25] which in principle can also be used to generate order-by-order expansions of the Matsubara correlator [28]. However, the high expansion orders achieved by Dyn-HTE in this work have not been matched by any other approach.
One experimental observable directly available from Dyn-HTE is the static susceptibility for arbitrary site-pairs. We consider the AFM $S=1/2$ Heisenberg chain and compare the static susceptibility from Dyn-HTE against error controlled quantum Monte Carlo (QMC). We also report the static susceptibility for the frustrated triangular lattice model. As an application we employ the static susceptibility of Dyn-HTE to showcase the accuracy of a simple approximate parametrization of its momentum dependence as suggested recently under the name renormalized mean-field form.
Finally, as motivated above, a main application of the Matsubara correlator is its analytical continuation to the real-frequency dynamical spin structure factor [12]. A large set of methods like QMC or pseudo-fermion based diagrammatic approaches [29, 30, 31, 32, 33] produce approximate numerical correlator data on a limited set of points on the Matsubara axis. In such a situation, despite recent advances [34, 35, 36, 37, 38], this analytical continuation is an ill-conditioned and error-prone procedure. For Dyn-HTE, in contrast, the obtained exact expansion can be regrouped in form of a high-frequency expansion (in inverse Matsubara frequency). In our companion work [39] we show that these high-frequency expansion coefficients can be identified with the frequency moments of the (real-frequency) spectral function from which the dynamical structure factor can be reconstructed by standard methods. This means that Dyn-HTE allows to bypass the standard ill-defined analytical continuation procedure.
@@ II Heisenberg spin model
We consider a system of length-$S$ quantum spins with operators $S^{\alpha}_{i}$ where $\alpha=x,y,z$. The subscript i=1,2,,Ni=1,2,...,N refers to the site at position $\mathbf{r}_{i}$ of an arbitrary lattice $\mathcal{L}$. The spins interact via Heisenberg exchange characterized by a single coupling constant $J$ along an arbitrary subset of all N(N1)/2N(N-1)/2 site pairs which we call bonds $(ii^{\prime})$. This includes the important case of symmetry-related nearest-neighbor interactions, but also all-to-all or spatially disordered (but equal) interactions like in a lattice with vacancies. The Hamiltonian reads
@@ III Matsubara spin correlator and Dyn-HTE
where $\nu_{m}=2\pi Tm$ with $m\in\mathbb{Z}$ is a (bosonic) Matsubara frequency and $S_{i}^{z}(\tau)=e^{H\tau}S_{i}^{z}e^{-H\tau}$ denotes imaginary time evolution. The operator $\mathcal{T}$ enforces imaginary time ordering, the operator with larger imaginary time argument is moved to the left. Note that the double integral is often replaced by a single integral over the difference $\tau-\tau^{\prime}$ but the form in (4) will prove useful later. Due to spin rotation symmetry in the Heisenberg Hamiltonian, Eq. (4) also determines correlations for all other spin-flavor combinations, $G^{\alpha\alpha^{\prime}}=\delta_{\alpha\alpha^{\prime}}G^{zz}$ and the $zz$ superscript of $G^{zz}$ is dropped in the following to ease notation. The spatial Fourier transform of Eq. (4) reads
@@
As we show in Sec. IV, the Dyn-HTE of the Matsubara correlator (4) takes the form of a double expansion in $x=J/T$ and inverse Matsubara frequency $1/\nu_{m}$. To order $n_{max}$ in $x$, Dyn-HTE reads
@@
$\begin{cases}p_{ii^{\prime}}^{(0)}(x)\!\!&\!:\!m=0\\ \sum_{r=1}^{r_{max}}p_{ii^{\prime}}^{(2r)}(x)\left(\frac{x}{2\pi m}\right)^{2r}\!\!&\!:\!m\neq 0\end{cases}\!+O(x^{n_{max}+1}),$ (7)
where $r_{max}=\left\lfloor n_{max}/2\right\rfloor$ is the integer floor of $n_{max}/2$. The dimensionless polynomials $p_{ii^{\prime}}^{(2r)}(x)$ for are of degree $n_{max}-2r$ with (real) rational coefficients, here $r=0,1,2,$.... Depending on the lattice and site-pair $ii^{\prime}$, some of the polynomial coefficients can be zero. Note that the form (7) is consistent with the symmetries of the Matsubara correlator discussed above.
Our open-source numerical implementation [21] provides the exact polynomial coefficients in Eq. (7) as rational numbers for arbitrary lattices $\mathcal{L}$, all site-pairs $ii^{\prime}$ therein. Currently spin lengths $S\in\{1/2,1\}$ and maximum expansion order $n_{max}=12$ are available.
@@
At finite frequency $m\!\neq\!0$, the Matsubara correlator does not directly represent an observable physical quantity. It is mainly considered for the simplicity of the diagrammatic field-theoretical framework which can be employed for its computation [12]. Usually, numerical data for the Matsubara correlator $G_{ii^{\prime}}(i\nu_{m})$ evaluated at a finite set of frequencies $\nu_{m}$ is linked to its real-frequency version by analytic continuation as discussed in Sec. I. Dyn-HTE allows to take a different and more stable route and never evaluates numerical values for $G_{ii^{\prime}}(i\nu_{m})$: As we show in our companion letter [39], the high-frequency expansion coefficients from Eq. (7) are in direct correspondence to the short-time expansion coefficients of the real-time correlator,
$p_{ii^{\prime}}^{(2r)}\sim\partial_{t}^{2r-1}\!\!\left\langle S_{i}^{z}(t)S_{i^{\prime}}^{z}\right\rangle\!|_{t=0}.$ (9)
The object on the right-hand side of Eq. (9) can also be identified with the (real-frequency) moments of the spectral function [41]. Using equations of motion, the moments can be unraveled as equal-time correlators of 2r12r-1 fold nested commutators, $\sim\!\left\langle[...[S_{i}^{z},\!H],\!H],...,\!H]S_{i^{\prime}}^{z}\right\rangle$. Even for moderate $r$, these are hard to compute by standard means beyond the case of one dimension and $T=\infty$. This explains the value of Dyn-HTE’s access to (the HTE of) $p_{ii^{\prime}}^{(2r)}$ for these cases. The full spectral function and dynamical structure factor can be reconstructed based on a moderate number of frequency moments via continued fraction representations [42, 41, 43, 44, 45]. In our companion letter [39] we discuss this in detail and show benchmarks and applications centered around the dynamical structure factor. In contrast, the remainder of the current work will be concerned with the efficient evaluation of the Dyn-HTE, i.e. the polynomials $p_{ii^{\prime}}^{(2r)}(x)$ in Eq. (7). Here we also present, benchmark and discuss results for the Matsubara correlator in imaginary frequency.
@@ IV Derivation of Dyn-HTE
In this central technical section we derive the Dyn-HTE as in Eq. (7) and provide an efficient algorithm for its calculation. To start, we split the Hamiltonian (1) into a non-interacting part $H_{0}=0$ (vanishing in the absence of a magnetic field) and an interacting part $V$ (here the full $H$) and define the dimensionless expansion coefficients $c_{ii^{\prime}}^{(n)}(i\nu_{m})$ of the Matsubara correlator via
@@
The lowest-order $n=0$ term represents the Curie susceptibility of a free spin which is local and static (non-zero only for $m=0$), $c_{ii^{\prime}}^{(0)}(i\nu_{m})=\delta_{ii^{\prime}}\delta_{0,m}S(S+1)/3$.
@@ IV.1 Expansion coefficients
General perturbation theory for the Matsubara correlator [12] provides an expression for the remaining expansion coefficients in Eq. (10) at order $n=1,2,$...
$c_{ii^{\prime}}^{(n)}(i\nu_{m})=\frac{T^{n+2}}{n!}\sum_{b_{1},...,b_{n}}\int_{0}^{\beta}e^{i\nu_{m}(\tau-\tau^{\prime})}\mathrm{d}\tau_{1}...\mathrm{d}\tau_{n}\mathrm{d}\tau\mathrm{d}\tau^{\prime}\left\langle\mathcal{T}V_{b_{1}}(\tau_{1})...V_{b_{n}}(\tau_{n})S_{i}^{z}(\tau)S_{i^{\prime}}^{z}(\tau^{\prime})\right\rangle_{0}^{\mathrm{V-con}}.$ (11)
The sum is over bonds $b_{l}=(i_{l}i_{l}^{\prime})$ and operator time arguments as in $V_{b_{1}}(\tau_{1})$ now refer to (interaction-picture) evolution with respect to $H_{0}$ only. This is trivial for our case, $H_{0}=0$, e.g. $V_{b_{1}}(\tau_{1})=V_{b_{1}}$. Nevertheless we must keep the time arguments because by virtue of $\mathcal{T}$ they determine the ordering of the string of generally non-commuting operators. These operators are thus distinguishable regardless of the chosen bonds $b_{l}$. The $\left\langle...\right\rangle_{0}$ denotes a thermal average with respect to Hamiltonian $H_{0}$. Since $H_{0}=0$, the former is simply an (effectively) infinite temperature average or normalized trace over the whole $N$-site spin Hilbert-space, $\left\langle...\right\rangle_{0}=\mathrm{tr}[...]/(2S+1)^{N}$ which factorizes according to site-index, e.g. $\left\langle S_{1}^{z}S_{2}^{+}S_{1}^{z}S_{2}^{-}\right\rangle_{0}=\left\langle S_{1}^{z}S_{1}^{z}\right\rangle_{0}\left\langle S_{2}^{+}S_{2}^{-}\right\rangle_{0}$ with the on-site operator order maintained.
The superscript $\mathrm{V-con}$” on the right-hand side of Eq. (11) refers to the $V$-connected correlator which contains all contributions to order $V^{n}$, also those from the expansion of the partition function $Z$ in the denominator on the right-hand side of Eq. (2). Following Ref. 46, it can be written via recursive subtractions from the ”full” correlator [without $\mathrm{V-con}$, c.f. (2)],
@@
[…]
@@
$c_{ii^{\prime}}^{(n)}(i\nu_{m})$ $=$ $\sum_{b_{1},...,b_{n}}\frac{T^{n+2}}{n!}\int_{0}^{\beta}e^{i\nu_{m}(\tau-\tau^{\prime})}\mathrm{d}\tau_{1}...\mathrm{d}\tau_{n}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
[…]